OpenStudy (anonymous):

21. Write the equation of a line that is perpendicular to the given line and that passes through the given point. y - 3 = 8/3(x + 2); (-2, 3)

OpenStudy (mertsj):

What is the given line?

OpenStudy (anonymous):

y-3=8/3(x+2) is the line given y=(8x+16)/3 +3 y=8/3x + 16/3+9/3 y=8/3x+25/3 perpendicular has the opposite slope. i.e. m1.m2=-1 if our slope(m1) is 8/3, the 8/3.m2=-1 or 8.m2=-3, m2=-3/8. check: m1.m2=-1, 8/3 x -3/8 = -24/24 = -1 correct. so the new slope will be -3/8 using our equation y-y1=m(x-x1) and our point (-2,3) we find the new line: y-3=-3/8(x-[-2]) y-3=-3(x+2)/8 y=-3x/8-6/8 simplify: y=(-3x-6)/8 thats your answer

OpenStudy (anonymous):

thank you that helps